OpenStudy (anonymous):

MEDAL & FAN ANYBODY ? If a cone has a volume of 18π cubic feet, and the height is 4 feet, what is the radius? A. square root 9.5 ft B. square root 11.5 ft C. square root 13.5 ft D. square root 15.5 ft

OpenStudy (anonymous):

V = pi*r^2 *(h/3) \[r = \sqrt{3*\frac{ v }{ \pi*h }}\]

OpenStudy (anonymous):

Put in 18pi for v and 4 for h and you get 3.67

OpenStudy (anonymous):

or the square root of 13.5